Express the concentration of solution in % w/v containing 12 g sodium chloride and 50 ml of water.

1
Expert's answer
2022-05-23T02:19:03-0400

1. C% = mNaCl / mtotal * 100 = 12 / (12+50) * 100 = 19.35 %

2. CM = n / V

n = mNaCl / MNaCl = 12 / 40 = 0,3 mol

CM = 0,3mol / 0,05 litres = 6 mol/l

ANSWER: 19,35 % NaCl of solution;

6 mol/l NaCl of solution.